OpenStudy (anonymous):

solve: log base 2 √5

OpenStudy (anonymous):

\[\log_{2}\sqrt{5}=\log_{2}{5^\frac{1}{2}} = \frac{1}{2} \log_{2}5=\frac{\log 5 }{2\log2} \]